Solve:
One number is 11 less than a second number. Twice the second number is 52 more than 4 times the first. Find the two numbers.
: Solve:
One number is 11 less than a second number. Twice the second number is 52 more than 4 times the first. Find the two numbers.

Answer by 24HoursTutor.com(40)

Let the two numbers be x and y. Now, we understand from the first sentence that the first number which we have assumed to be x is 11 less than the second number y. Which means that if we add eleven to the first number we will get the second number. Thus, we get the equation as :
x + 11 = y
Now, the second sentence tells us that twice the second number, y in our case is equal to 4 times the first number(x) and 52 more. So we have :
2y = 52 + 4x
Since we have an equation for y , we substitute it in this equation to get :
2(x+11) = 52 + 4x
2x + 22 = 52 + 4x
2x - 4x = 52 -22
- 2x = 30
x = 30 / -2
x = -15
Now, since we have the value of x, we can get the value of y by :
y = x + 11
y = -15 + 11
y = -4
Ans.: The first number is -15 and the second number is -4. You can cross check your answer by substituting the value in both the equations and seeing if they still hold true.
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr